OpenStudy (amistre64):

if \(\phi): RxR \(\to\) R, and ker\(\phi\)={(x,y)\(\in\) RxR | x+y=0}; what is RxR/ker\(\phi\) isomorphic to. What is a description of that?

OpenStudy (amistre64):

almost got the coding right lol ....

OpenStudy (amistre64):

\(\phi\): RxR \(\to\) R

OpenStudy (amistre64):

i spose my real question is, what does a factor group describe? my current thinking is that a factor group is a coset, H, that divvys up the group into equal partitions. Does H have to be normal in the group? I think ker\(\phi\) is normal to the group ... last night i was thinking maybe the desciption is: since the kernel of phi is the set of ordered pairs (x,y) such that y=-x, this is a coset of RxR such that it is divided into partitons of parallel lines such that form some (a,b)H is the set of ordered pairs (x,y) such that y-b = -(x-a)

OpenStudy (anonymous):

Just curious to which field of mathematics your question belongs to ?

OpenStudy (amistre64):

it was a question on my abstract algebra final

OpenStudy (amistre64):

so im assuming none of the fields :)

OpenStudy (anonymous):

Is \(\mathbb{R}\times \mathbb{R}\) being considered as a group under addition?

OpenStudy (amistre64):

dunno ... i think the definition it used was:\[\phi(\,(x,y)\,)=x+y\]

OpenStudy (amistre64):

and stated phi was a homomorphism

OpenStudy (amistre64):

addition seems familiar .... might have been addition of vectors in RxR

OpenStudy (anonymous):

ok but in order for it to make sense, then RxR/ker assumes that RxR is something a group, ring, something

OpenStudy (amistre64):

addition makes the most sense since we were spose to show that:\[\phi(ab)=\phi(a)\phi(b)\]

OpenStudy (amistre64):

then lets go with group, RxR with operation of addition of vectors

OpenStudy (anonymous):

not that i know the answer, but i have a way to think about it, informally at least things in ker f are \((a,-a)\) so for example \((5,2)\) would be in the same coset as \((3,0)\) because if you are thinking of it as a group under addition, then \((5,2)=(3,0)+(2,-2)\)

OpenStudy (anonymous):

well i screwed that up, didn't i?

OpenStudy (anonymous):

what i meant was the same coset as \((7,0)\)

OpenStudy (anonymous):

in other words, \((b,a)\) would be in the same coset as \((b+a,0)\)

OpenStudy (anonymous):

making it isomorphic, if my intuition is correct, to \(\mathbb{R}\)

OpenStudy (anonymous):

i would only bet $6 on this answer, because i am only 60% sure it is right

OpenStudy (amistre64):

thats making sense, since the coset is made of the elements in RxR, then aH or Ha is the operation in RxR

OpenStudy (anonymous):

unfortunately i gotta run to class i will think about it more later but yes, what you said, except since the operation is (i assume) pointwise addition, it would make more sense to write \(a+H\) where of course \(a=(x,y)\)

OpenStudy (anonymous):

another intuitive way to think about it RxR is a plane x + y = 0 is a line mod out by the line, you get a line

OpenStudy (anonymous):

or yet another way all other cosets are parallel to \(x+y=0\) they are all the same, just differ by the y intercept boy am i going to be embarrassed if this is wrong ....

OpenStudy (amistre64):

the parallel line was what i was thinking, similar to a notion of Z4 forms equivalence classes by considering all the remaniders left when dividing by 4; this current setup may just as well be a set of equivalence classes of affine lines by considereing their y intercepts.